# Perfect Edge Domination in $P_{6}$-free Graphs and in Graphs Without Efficient Edge Dominating Sets

### v1

@@ 4.2.1 $V(K)\cap B=\emptyset$
Let $Z$ denote the set of vertices with degree zero in $G_{0}$ and no neighbor in $S_{1}\cup X_{1}$. Suppose $\pi$ is a valid $3$-coloring of $G$, and let $S_{2}^{\pi}$ (resp. $X_{2}^{\pi}$) be the subset of vertices $v$ in $S_{2}$ (resp. $X_{2}$) such that $\pi(u)=w$ for every $u\in N_{G_{0}}(v)$ with $d_{G_{0}}(u)\geq 1$. We define $H_{\pi}$ as the induced bipartite subgraph $G[S_{2}^{\pi}\cup X_{2}^{\pi}\cup Z]$ of $G$. The following lemma addresses the trivial connected components of $G_{0}$.
@@
Let $G$ be a connected $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$, with $R\subseteq W$ and $S\subseteq Y$. Then, $G$ admits a valid $3$-coloring $\pi$ such that each connected component $C$ of $H_{\pi}$ has a valid $3$-coloring in which the vertices of $V(C)\cap S_{2}^{\pi}$ are assigned the yellow color. Moreover:
@@
    Such a $3$-coloring of $C$ is valid if and only if $V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})$ can be partitioned into neighborhoods of non-white vertices in $V(C)\cap Z$, and for each $z\in V(C)\cap Z$, the vertex $z$ is yellow if and only if $|N_{C}(z)|=1$. Hence, any PED-set of $C$ contains exactly $|V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})|$ edges.
@@
    Some vertex $z$ of $V(C)\cap Z$ has degree $|V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})|$ in $C$. A valid $3$-coloring can be obtained by assigning the white color to all vertices of $V(C)\cap Z$ except $z$.
@@
    Other valid $3$-colorings can be obtained by assigning the white color to $z$ and considering the valid 3-colorings of $C-\{z\}$ (note that $C-\{z\}$ may be disconnected).
@@
Let $\pi$ be a valid $3$-coloring for $G$. Hence, $H_{\pi}$ has a valid $3$-coloring $\pi^{\prime}$ defined as $\pi$ restricted to $V(H_{\pi})$; i.e, $\pi^{\prime}(v)=\pi(v)$ for every vertex $v\in V(H_{\pi}$). By Lemma 4.2, $H_{\pi}$ has a vertex $z^{*}\in Z$ adjacent to each vertex of $S_{2}^{\pi}$. Therefore, if $\pi$ is an optimum valid $3$-coloring of $G$, $\pi^{*}$ defined as follows: $\pi^{*}(u)=\pi(u)$, for each $u\in G-Z$, $\pi^{*}(z^{*})=b$ whenever $|N(z^{*})\cap S_{2}^{\prime}|\geq 2$, $\pi^{*}(z^{*})=y$ whenever $|N(z^{*})\cap S_{2}^{\prime}|=1$, and $\pi^{*}(z)=w$ for every $z\in Z\setminus\{z^{*}\}$ turns into a valid $3$-coloring for $G$.
Statement 4. follows from the fact that if for a vertex $v$ is assigned white in valid $3$-coloring, the $3$-coloring $\pi$ restricted to $V(G-\{v\})$ is also valid. ∎
@@ 4.2.2 $R\cap B\neq\emptyset$ and $S\cap B=\emptyset$
Now, we will examine $P_{6}$-free graphs having a valid $3$-coloring and a dominating (not necessarily induced) complete bipartite subgraph $K$ with at least one black vertex.
We begin by analyzing the valid partial $3$-coloring of $K$ in which exactly one of the sets of the bipartition contains black vertices.
@@
Let $G$ be a graph with a valid $3$-coloring $(B,Y,W)$. Suppose that $G$ has a dominating (not necessarily induced) complete bipartite subgraph $K$ with a bipartition $\{R,S\}$ such that $R\cap B=\{r\}$ and $S\subseteq Y$. Then the following conditions hold:
@@
The first three statements follow directly from the definition of a valid $3$-coloring.
Since each vertex $x\in N$ is adjacent to at least one vertex in $R$, it follows from Statement 2 that $x$ is a non-white vertex. If $d^{N}(x)\geq 2$, then $x\in B$, and $N_{K}(x)=\{r\}$. If $d^{N}(x)=0$, since $N_{G}(x)\cap V(K)\subseteq R$, then $x\in Y$ and must be adjacent to $r$. Now suppose that $d^{N}(x)=1$ and $x$ has at least two neighbors in $R\cup M$. Then $x$ has at least one white neighbor in $R\cup M$, and one non-white neighbor in $N$, which implies that $x$ must be a yellow. In particular, $x$ cannot be adjacent to $r$. ∎
@@
It is always advisable to choose the one that has the lowest degree in $G$ to obtain a smaller PED-set. Once $r$ is chosen, the valid 3-coloring can be completed as follows: the black vertices of $N$ are exactly neighbors of $r$ and have degree at least one in $G[N]$.
@@ 4.2.3 $R\cap B\neq\emptyset$ and $S\cap B\neq\emptyset$
Now, let us consider $P_{6}$-free graphs having a dominating complete bipartite subgraph $K$ with a partial $3$-coloring where both sets of the bipartition contain black vertices. It is important to note that in this case $V(K)$ does not have any white vertices.
@@
Let $G$ be a connected $P_{6}$-free graph with a valid $3$-coloring $(B,Y,W)$. If $G$ has a dominating complete bipartite subgraph $K$ such that $R\cap B\neq\emptyset$ and $S\cap B\neq\emptyset$, then one of the following conditions holds:
@@
Let $G$ be a connected $P_{6}$-free graph with a maximal dominating complete bipartite subgraph $K$, having a bipartition $\{R=\{r\},S=N(r)\}$, and let $t=|S|$. Suppose that $G$ admits a valid $3$-coloring $\pi$ such that $\pi(r)=b$ and $S\cap B\neq\emptyset$. Then the following conditions hold:
@@ 5 Unweighted minimum PED
    For each possible partial $3$-coloring of $C$ of types (b), (c) and (e), attempt to extend it to a total $3$-coloring of $G$ that corresponds to a proper PED-set according to Theorem 4.2. If the extension is valid, compare it with the best current PED-set $\mathcal{P}$ and update $\mathcal{P}$ if the new one is better. Validity checking of a total $3$-coloring requires $O(m+n)$-time.
    The number of possible total $3$-colorings is $O(1)$ except those of type (e). Since the number of possible colorings for each type (b) and (c) is $6$ (see Lemma 4.1 and Theorem 4.2), each extension can be done in $O(m+n)$-time. Therefore, the generation and validation of these 12 total $3$-colorings can clearly be done in overall $O(m+n)$-time.
    Next, we analyze partial $3$-colorings of type (e) for $C$. There are exactly two such colorings. For each of them, one of the following situations occurs (see Theorem 4.2):
@@
      No valid total $3$-coloring can be extended.
@@
      Exactly one valid total $3$-coloring can be extended, which corresponds to a DIM (i.e., $G[Y]$ is $1$-regular).
@@
      Multiple total $3$-colorings can be extended ($G[Y]$ is not $1$-regular). If one of them is valid, then all of them are valid and correspond to a proper PED-sets. These PED-sets have the same number of edges, since the corresponding valid $3$-colorings share the same set of yellow vertices and exactly one black vertex with a specific degree. In this case, it suffices to consider any one of them.
    Again, at most two significant total $3$-colorings (corresponding to type (e)) are generated and checked in $O(n+m)$ time.
@@
      Configuration I: $K$ has a partial $3$-coloring with no black vertices. Perform the non-black-vertex sub-algorithm. See Subsection 4.2.1.
@@
The list $\Pi$ contains valid partial $3$-colorings found so far. At various stages, initial valid partial $3$-colorings are generated and stored in $\Pi$. Subsequent steps attempt to extend these partial $3$-colorings into valid total $3$-colorings of $G$, or determines that such an extension is not possible.
@@
    Assign white color to the vertices of $R$, and a yellow color to the vertices of $S$. Then, invert the coloring of each set in the bipartition (see Lemma 4.5), and repeat the following substeps for each of these two initial valid partial $3$-colorings:
@@
      Check that $X\cup S$ is a dissociation set: compute $G[X\cup S]$, $S_{1}:=\{s\in S\mid d_{G[X\cup S]}(s)=1\}$, $S_{2}:=\{s\in S\mid d_{G[X\cup S]}(s)=0\}$, $X_{1}:=\{x\in X\mid d_{G[X\cup S]}(x)=1\}$ and $X_{2}:=\{x\in X\mid d_{G[X\cup S]}(x)=0\}$. If it fails (that is, $S\neq S_{1}\cup S_{2}$ or $X\neq X_{1}\cup X_{2}$), then skip to the next initial valid partial $3$-coloring of the step (this one cannot be extended to a valid total $3$-coloring of $G$).
@@
      If a vertex $v\in S_{2}\cup X_{2}$ has $d^{V(G_{0})}_{G}(v)=0$, then skip to the next initial valid partial $3$-coloring of the step (this one cannot be extended to a valid total $3$-coloring of $G$). See Remark 4.3.
@@
      For each $v\in S_{1}\cup X_{1}$, assign white color to each vertex $u\in N_{G}(v)\cap V(G_{0})$, and assign a white/yellow coloring to the bipartite connected component $C(u)$ of $u$ in $G_{0}$, compatible with $u$, whenever $u$ is uncolored. If some $u\in N(v)\cap V(G_{0})$ was previously assigned a non-white color, skip to the next initial valid partial $3$-coloring of the step (this one cannot be extended to a valid total $3$-coloring of $G$).
@@
      If all connected components of $\mathcal{C}$ are now colored, store the partial $3$-coloring in $\Pi$. Otherwise, let $C\in\mathcal{C}$ be an uncolored component of type I. In this case, there exists some vertex $v\in S_{2}$ that has a neighbor $c$ and a non-neighbor $d$ in $C$, such that $cd$ is an edge of $C$. If $d$ has a exactly one neighbor $v^{\prime}$ in $S_{2}$, and $v^{\prime}$ is adjacent to $c$, then assign yellow to each vertex in the partition class of $C$ containing $d$, and white to each vertex in the other class ($c$ has 2 yellow neighbors $v$ and $v^{\prime}$ which implies that $c\notin Y$, as $\{v^{\prime},c,d\}$ forms a triangle which implies that $c\notin B$, $c\in W$ and $d\in Y$). Repeat this procedure as long as components satisfying this condition remain. Otherwise, check the resulting partial $3$-coloring is valid (that is any yellow vertex has at most one colored non-white neighbor and the colored non-yellow vertices form an independent set)? If the check fails, then skip to the next initial valid partial $3$-coloring of the step (this one cannot be extended to a valid total $3$-coloring of $G$). Otherwise, consider up to three extensions of the current partial $3$-coloring as follows:
@@
        Assign black to all vertices of $C$. If there is some vertex $z\in S_{2}\cup X_{2}$ with $d^{V(C)}_{G}(z)\geq 2$ then skip to the next alternative (that partial $3$-coloring cannot be extended to a valid $3$-coloring of $G$). Otherwise, for each $z\in S_{2}\cup X_{2}$ with $d^{V(C)}_{G}(z)=1$, assign white color to each $u\in N(z)\cap(V(G_{0})\setminus V(C))$, and assign a compatible white/yellow coloring to $C(u)$ compatible with $u$, whenever $u$ is uncolored. If any such $u$ was previously colored non-white (that partial $3$-coloring cannot be extended to a valid $3$-coloring of $G$) skip to the next alternative. Otherwise, store it in $\Pi$.
@@
        If $c$ has exactly one neighbor in $S_{2}$, do the following: assign yellow to $c$ and assign a compatible white/yellow coloring to $C(c)$ compatible with $c$. Then assign white to each vertex $y\in N(v)\cap(V(G_{0})\setminus\{c\})$, and assign a white/yellow coloring to $C(y)$ whenever $y$ is uncolored. If any such $y$ was previously colored with a non-white color (that partial $3$-coloring cannot be extended to a valid $3$-coloring of $G$) skip to the next alternative. Otherwise, store the resulting partial $3$-coloring in $\Pi$.
@@
        If $d$ has exactly one neighbor $v^{\prime}\in S_{2}$ (clearly, $c\notin N_{G}(v^{\prime})$ and $v^{\prime}$ is another indicator of $C$), do the following: assign yellow to $d$ and a compatible white/yellow coloring to $C(d)$ compatible with $d$. Then assign white to each vertex $y\in N(v^{\prime})\cap(V(G_{0})\setminus\{d\})$, and assign a compatible white/yellow coloring to $C(y)$ whenever $y$ is uncolored. If any such $y$ was previously assigned a non-white color (that partial $3$-coloring cannot be extended to a valid $3$-coloring of $G$) skip to the next initial valid partial $3$-coloring. Otherwise, store that partial $3$-coloring in $\Pi$.
    The time complexity of this step is $O(m+n)$. The list $\Pi$ has at most 6 partial $3$-colorings.
@@
    For each valid partial $3$-coloring of $\Pi$, repeat the following substeps.
@@
        Otherwise, discard that partial $3$-coloring.
@@
      Compute $S^{\pi}_{2}$, $X^{\pi}_{2}$ and $Z$, where $\pi$ is the resulting partial $3$-coloring of (a). Consider the graph $H_{\pi}$. For each connected component $C$ of $H_{\pi}$ that contains some vertex in $S^{\pi}_{2}\cup X^{\pi}_{2}$, find a vertex $v$ that is adjacent to every vertex in $V(C)\cap(S^{\pi}_{2}\cup X^{\pi}_{2})$. Assign the yellow color to $v$ if $d_{C}(v)=1$, or the black color if $d_{C}(v)>1$, and assign the white color to every other uncolored vertex in that component. If no such a vertex exists for some connected component, discard that partial $3$-coloring (see Lemma 4.8).
@@
      If the resulting $3$-coloring is a valid total $3$-coloring, compare it with the current best PED-set $\mathcal{P}$ and update $\mathcal{P}$ if it is better.
@@
    Check that $R$ and $S$ are both independent sets. If it fails, then it is not possible to extend to a valid $3$-coloring.
@@
    $M=\{v\in V(G)\setminus V(K):N^{S}_{G}(v)\neq\emptyset\}$ and check that $R\cup M$ is an independent set. Again, if it fails, then it is not possible to extend to a valid $3$-coloring.
@@
    Any vertex $r^{\prime}$ of $R$ that have vertices of (a) and (b) as neighbors and vertices of (c) as non-neighbors can be selected as $r$ and complete a valid $3$-coloring as follow: $r\in B$, $R\setminus\{r\}\subseteq W$, and any neighbor (non-neighbor) $x$ of $r$ with $d^{N}(x)=1$ must be black (yellow). It is always advisable to choose $r^{\prime}$ with lowest degree in $G$ as $r$ to obtain a smaller PED-set.
@@
      Any of these combinations can complete a valid $3$-coloring: assign either yellow or black to each vertex in $X$, depending on whether it has exactly degree one in GWG-W or not. The corresponding PED-set is non-trivial if and only if some connected component $C$ has taken (ii) as option. For unweighted version problem, it is always convenient to choose option (ii) for every connected component. Consequently, $S_{2}^{\prime}\subseteq Y$ and $Z\setminus X\subseteq W$.
@@
      If $C$ is non-bipartite, or if both $c$ and $d$ have at least two neighbors in $S$, then stop. In this case, the only valid total $3$-coloring corresponds to the trivial PED-set (Lemma 4.11, item 4).
@@
      Otherwise, $C$ is bipartite and explore at most two possible extensions of the partial $3$-coloring:
@@
        • If the resulting $3$-coloring is valid, compare it with the current best PED-set $\mathcal{P}$ and update if necessary.
@@
        • If the resulting $3$-coloring is a valid, compare it with the current best PED-set $\mathcal{P}$ and update if necessary.

### v2

@@ 4.2.1 $V(K)\cap B=\emptyset$
Let $Z$ denote the set of vertices with degree zero in $G_{0}$ and no neighbor in $S_{1}\cup X_{1}$. Suppose $\pi$ is a valid 33-coloring of $G$, and let $S_{2}^{\pi}$ (resp. $X_{2}^{\pi}$) be the subset of vertices $v$ in $S_{2}$ (resp. $X_{2}$) such that $\pi(u)=w$ for every $u\in N_{G_{0}}(v)$ with $d_{G_{0}}(u)\geq 1$. We define $H_{\pi}$ as the induced bipartite subgraph $G[S_{2}^{\pi}\cup X_{2}^{\pi}\cup Z]$ of $G$. The following lemma addresses the trivial connected components of $G_{0}$.
@@
Let $G$ be a connected $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$, with $R\subseteq W$ and $S\subseteq Y$. Then, $G$ admits a valid 33-coloring $\pi$ such that each connected component $C$ of $H_{\pi}$ has a valid 33-coloring in which the vertices of $V(C)\cap S_{2}^{\pi}$ are assigned the yellow color. Moreover:
@@
    Such a 33-coloring of $C$ is valid if and only if $V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})$ can be partitioned into neighborhoods of non-white vertices in $V(C)\cap Z$, and for each $z\in V(C)\cap Z$, the vertex $z$ is yellow if and only if $|N_{C}(z)|=1$. Hence, any PED-set of $C$ contains exactly $|V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})|$ edges.
@@
    Some vertex $z$ of $V(C)\cap Z$ has degree $|V(C)\cap(S_{2}^{\pi}\cup X_{2}^{\pi})|$ in $C$. A valid 33-coloring can be obtained by assigning the white color to all vertices of $V(C)\cap Z$ except $z$.
@@
    Other valid 33-colorings can be obtained by assigning the white color to $z$ and considering the valid 3-colorings of $C-\{z\}$ (note that $C-\{z\}$ may be disconnected).
@@
Let $\pi$ be a valid 33-coloring for $G$. Hence, $H_{\pi}$ has a valid 33-coloring $\pi^{\prime}$ defined as $\pi$ restricted to $V(H_{\pi})$; i.e, $\pi^{\prime}(v)=\pi(v)$ for every vertex $v\in V(H_{\pi}$). By Lemma 4.2, $H_{\pi}$ has a vertex $z^{*}\in Z$ adjacent to each vertex of $S_{2}^{\pi}$. Therefore, if $\pi$ is an optimum valid 33-coloring of $G$, $\pi^{*}$ defined as follows: $\pi^{*}(u)=\pi(u)$, for each $u\in G-Z$, $\pi^{*}(z^{*})=b$ whenever $|N(z^{*})\cap S_{2}^{\prime}|\geq 2$, $\pi^{*}(z^{*})=y$ whenever $|N(z^{*})\cap S_{2}^{\prime}|=1$, and $\pi^{*}(z)=w$ for every $z\in Z\setminus\{z^{*}\}$ turns into a valid 33-coloring for $G$.
Statement 4. follows from the fact that if for a vertex $v$ is assigned white in valid 33-coloring, the 33-coloring $\pi$ restricted to $V(G-\{v\})$ is also valid. ∎
@@ 4.2.2 $R\cap B\neq\emptyset$ and $S\cap B=\emptyset$
Now, we will examine $P_{6}$-free graphs having a valid 33-coloring and a dominating (not necessarily induced) complete bipartite subgraph $K$ with at least one black vertex.
We begin by analyzing the valid partial 33-coloring of $K$ in which exactly one of the sets of the bipartition contains black vertices.
@@
Let $G$ be a graph with a valid 33-coloring $(B,Y,W)$. Suppose that $G$ has a dominating (not necessarily induced) complete bipartite subgraph $K$ with a bipartition $\{R,S\}$ such that $R\cap B=\{r\}$ and $S\subseteq Y$. Then the following conditions hold:
@@
The first three statements follow directly from the definition of a valid 33-coloring.
Since each vertex $x\in N$ is adjacent to at least one vertex in $R$, it follows from Statement 2 that $x$ is a non-white vertex. If $d^{N}(x)\geq 2$, then $x\in B$, and $N_{K}(x)=\{r\}$. If $d^{N}(x)=0$, since $N_{G}(x)\cap V(K)\subseteq R$, then $x\in Y$ and must be adjacent to $r$. Now suppose that $d^{N}(x)=1$ and $x$ has at least two neighbors in $R\cup M$. Then $x$ has at least one white neighbor in $R\cup M$, and one non-white neighbor in NN, which implies that $x$ must be a yellow. In particular, $x$ cannot be adjacent to $r$. ∎
@@
It is always advisable to choose the one that has the lowest degree in $G$ to obtain a smaller PED-set. Once $r$ is chosen, the valid 3-coloring can be completed as follows: the black vertices of NN are exactly neighbors of $r$ and have degree at least one in $G[N]$.
@@ 4.2.3 $R\cap B\neq\emptyset$ and $S\cap B\neq\emptyset$
Now, let us consider $P_{6}$-free graphs having a dominating complete bipartite subgraph $K$ with a partial 33-coloring where both sets of the bipartition contain black vertices. It is important to note that in this case $V(K)$ does not have any white vertices.
@@
Let $G$ be a connected $P_{6}$-free graph with a valid 33-coloring $(B,Y,W)$. If $G$ has a dominating complete bipartite subgraph $K$ such that $R\cap B\neq\emptyset$ and $S\cap B\neq\emptyset$, then one of the following conditions holds:
@@
Let $G$ be a connected $P_{6}$-free graph with a maximal dominating complete bipartite subgraph $K$, having a bipartition $\{R=\{r\},S=N(r)\}$, and let $t=|S|$. Suppose that $G$ admits a valid 33-coloring $\pi$ such that $\pi(r)=b$ and $S\cap B\neq\emptyset$. Then the following conditions hold:
@@ 5 Unweighted minimum PED
    For each possible partial 33-coloring of $C$ of types (b), (c) and (e), attempt to extend it to a total 33-coloring of $G$ that corresponds to a proper PED-set according to Theorem 4.2. If the extension is valid, compare it with the best current PED-set $\mathcal{P}$ and update $\mathcal{P}$ if the new one is better. Validity checking of a total 33-coloring requires $O(m+n)$-time.
    The number of possible total 33-colorings is $O(1)$ except those of type (e). Since the number of possible colorings for each type (b) and (c) is 66 (see Lemma 4.1 and Theorem 4.2), each extension can be done in $O(m+n)$-time. Therefore, the generation and validation of these 12 total 33-colorings can clearly be done in overall $O(m+n)$-time.
    Next, we analyze partial 33-colorings of type (e) for $C$. There are exactly two such colorings. For each of them, one of the following situations occurs (see Theorem 4.2):
@@
      No valid total 33-coloring can be extended.
@@
      Exactly one valid total 33-coloring can be extended, which corresponds to a DIM (i.e., $G[Y]$ is 11-regular).
@@
      Multiple total 33-colorings can be extended ($G[Y]$ is not 11-regular). If one of them is valid, then all of them are valid and correspond to a proper PED-sets. These PED-sets have the same number of edges, since the corresponding valid 33-colorings share the same set of yellow vertices and exactly one black vertex with a specific degree. In this case, it suffices to consider any one of them.
    Again, at most two significant total 33-colorings (corresponding to type (e)) are generated and checked in $O(n+m)$ time.
@@
      Configuration I: $K$ has a partial 33-coloring with no black vertices. Perform the non-black-vertex sub-algorithm. See Subsection 4.2.1.
@@
The list $\Pi$ contains valid partial 33-colorings found so far. At various stages, initial valid partial 33-colorings are generated and stored in $\Pi$. Subsequent steps attempt to extend these partial 33-colorings into valid total 33-colorings of $G$, or determines that such an extension is not possible.
@@
    Assign white color to the vertices of $R$, and a yellow color to the vertices of $S$. Then, invert the coloring of each set in the bipartition (see Lemma 4.5), and repeat the following substeps for each of these two initial valid partial 33-colorings:
@@
      Check that $X\cup S$ is a dissociation set: compute $G[X\cup S]$, $S_{1}:=\{s\in S\mid d_{G[X\cup S]}(s)=1\}$, $S_{2}:=\{s\in S\mid d_{G[X\cup S]}(s)=0\}$, $X_{1}:=\{x\in X\mid d_{G[X\cup S]}(x)=1\}$ and $X_{2}:=\{x\in X\mid d_{G[X\cup S]}(x)=0\}$. If it fails (that is, $S\neq S_{1}\cup S_{2}$ or $X\neq X_{1}\cup X_{2}$), then skip to the next initial valid partial 33-coloring of the step (this one cannot be extended to a valid total 33-coloring of $G$).
@@
      If a vertex $v\in S_{2}\cup X_{2}$ has $d^{V(G_{0})}_{G}(v)=0$, then skip to the next initial valid partial 33-coloring of the step (this one cannot be extended to a valid total 33-coloring of $G$). See Remark 4.3.
@@
      For each $v\in S_{1}\cup X_{1}$, assign white color to each vertex $u\in N_{G}(v)\cap V(G_{0})$, and assign a white/yellow coloring to the bipartite connected component $C(u)$ of $u$ in $G_{0}$, compatible with $u$, whenever $u$ is uncolored. If some $u\in N(v)\cap V(G_{0})$ was previously assigned a non-white color, skip to the next initial valid partial 33-coloring of the step (this one cannot be extended to a valid total 33-coloring of $G$).
@@
      If all connected components of $\mathcal{C}$ are now colored, store the partial 33-coloring in $\Pi$. Otherwise, let $C\in\mathcal{C}$ be an uncolored component of type I. In this case, there exists some vertex $v\in S_{2}$ that has a neighbor $c$ and a non-neighbor $d$ in $C$, such that cdcd is an edge of $C$. If $d$ has a exactly one neighbor $v^{\prime}$ in $S_{2}$, and $v^{\prime}$ is adjacent to $c$, then assign yellow to each vertex in the partition class of $C$ containing $d$, and white to each vertex in the other class ($c$ has 2 yellow neighbors $v$ and $v^{\prime}$ which implies that $c\notin Y$, as $\{v^{\prime},c,d\}$ forms a triangle which implies that $c\notin B$, $c\in W$ and $d\in Y$). Repeat this procedure as long as components satisfying this condition remain. Otherwise, check the resulting partial 33-coloring is valid (that is any yellow vertex has at most one colored non-white neighbor and the colored non-yellow vertices form an independent set)? If the check fails, then skip to the next initial valid partial 33-coloring of the step (this one cannot be extended to a valid total 33-coloring of $G$). Otherwise, consider up to three extensions of the current partial 33-coloring as follows:
@@
        Assign black to all vertices of $C$. If there is some vertex $z\in S_{2}\cup X_{2}$ with $d^{V(C)}_{G}(z)\geq 2$ then skip to the next alternative (that partial 33-coloring cannot be extended to a valid 33-coloring of $G$). Otherwise, for each $z\in S_{2}\cup X_{2}$ with $d^{V(C)}_{G}(z)=1$, assign white color to each $u\in N(z)\cap(V(G_{0})\setminus V(C))$, and assign a compatible white/yellow coloring to $C(u)$ compatible with $u$, whenever $u$ is uncolored. If any such $u$ was previously colored non-white (that partial 33-coloring cannot be extended to a valid 33-coloring of $G$) skip to the next alternative. Otherwise, store it in $\Pi$.
@@
        If $c$ has exactly one neighbor in $S_{2}$, do the following: assign yellow to $c$ and assign a compatible white/yellow coloring to $C(c)$ compatible with $c$. Then assign white to each vertex $y\in N(v)\cap(V(G_{0})\setminus\{c\})$, and assign a white/yellow coloring to $C(y)$ whenever $y$ is uncolored. If any such $y$ was previously colored with a non-white color (that partial 33-coloring cannot be extended to a valid 33-coloring of $G$) skip to the next alternative. Otherwise, store the resulting partial 33-coloring in $\Pi$.
@@
        If $d$ has exactly one neighbor $v^{\prime}\in S_{2}$ (clearly, $c\notin N_{G}(v^{\prime})$ and $v^{\prime}$ is another indicator of $C$), do the following: assign yellow to $d$ and a compatible white/yellow coloring to $C(d)$ compatible with $d$. Then assign white to each vertex $y\in N(v^{\prime})\cap(V(G_{0})\setminus\{d\})$, and assign a compatible white/yellow coloring to $C(y)$ whenever $y$ is uncolored. If any such $y$ was previously assigned a non-white color (that partial 33-coloring cannot be extended to a valid 33-coloring of $G$) skip to the next initial valid partial 33-coloring. Otherwise, store that partial 33-coloring in $\Pi$.
    The time complexity of this step is $O(m+n)$. The list $\Pi$ has at most 6 partial 33-colorings.
@@
    For each valid partial 33-coloring of $\Pi$, repeat the following substeps.
@@
        Otherwise, discard that partial 33-coloring.
@@
      Compute $S^{\pi}_{2}$, $X^{\pi}_{2}$ and $Z$, where $\pi$ is the resulting partial 33-coloring of (a). Consider the graph $H_{\pi}$. For each connected component $C$ of $H_{\pi}$ that contains some vertex in $S^{\pi}_{2}\cup X^{\pi}_{2}$, find a vertex $v$ that is adjacent to every vertex in $V(C)\cap(S^{\pi}_{2}\cup X^{\pi}_{2})$. Assign the yellow color to $v$ if $d_{C}(v)=1$, or the black color if $d_{C}(v)>1$, and assign the white color to every other uncolored vertex in that component. If no such a vertex exists for some connected component, discard that partial 33-coloring (see Lemma 4.8).
@@
      If the resulting 33-coloring is a valid total 33-coloring, compare it with the current best PED-set $\mathcal{P}$ and update $\mathcal{P}$ if it is better.
@@
    Check that $R$ and $S$ are both independent sets. If it fails, then it is not possible to extend to a valid 33-coloring.
@@
    $M=\{v\in V(G)\setminus V(K):N^{S}_{G}(v)\neq\emptyset\}$ and check that $R\cup M$ is an independent set. Again, if it fails, then it is not possible to extend to a valid 33-coloring.
@@
    Any vertex $r^{\prime}$ of $R$ that have vertices of (a) and (b) as neighbors and vertices of (c) as non-neighbors can be selected as $r$ and complete a valid 33-coloring as follow: $r\in B$, $R\setminus\{r\}\subseteq W$, and any neighbor (non-neighbor) $x$ of $r$ with $d^{N}(x)=1$ must be black (yellow). It is always advisable to choose $r^{\prime}$ with lowest degree in $G$ as $r$ to obtain a smaller PED-set.
@@
      Any of these combinations can complete a valid 33-coloring: assign either yellow or black to each vertex in $X$, depending on whether it has exactly degree one in GWG-W or not. The corresponding PED-set is non-trivial if and only if some connected component $C$ has taken (ii) as option. For unweighted version problem, it is always convenient to choose option (ii) for every connected component. Consequently, $S_{2}^{\prime}\subseteq Y$ and $Z\setminus X\subseteq W$.
@@
      If $C$ is non-bipartite, or if both $c$ and $d$ have at least two neighbors in $S$, then stop. In this case, the only valid total 33-coloring corresponds to the trivial PED-set (Lemma 4.11, item 4).
@@
      Otherwise, $C$ is bipartite and explore at most two possible extensions of the partial 33-coloring:
@@
        • If the resulting 33-coloring is valid, compare it with the current best PED-set $\mathcal{P}$ and update if necessary.
@@
        • If the resulting 33-coloring is a valid, compare it with the current best PED-set $\mathcal{P}$ and update if necessary.